OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x − 1)? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 greater than or equal to x x less than or greater to 3

OpenStudy (ivettef365):

so you have 5x + 2 >= 7x + 2x -2

OpenStudy (ivettef365):

combine like terms

OpenStudy (anonymous):

I dont know how, I just started learning this stuff

OpenStudy (ivettef365):

ok, we can start by combining the 7x + 2x and we have 5x + 2 >= 9x -2

OpenStudy (ivettef365):

then you combine like terms by switching sides, change sign when switching numbers so you have 5x - 9x >= -2 -2

OpenStudy (ivettef365):

are you understanding so far ?

OpenStudy (texaschic101):

do you know what like terms are ?

OpenStudy (texaschic101):

we cant help you unless you respond